# Advancing Physiological Time Series Reconstruction and Imputation via Mixture of Receptive Fields and Experts Fusion

### v3

@@ 1 Introduction
Medical time series data have gained increasing attention with the advancement of deep neural networks. Among these, physiological time series signals, such as Electrocardiogram (ECG), Electroencephalogram (EEG), and Polysomnography (PSG), are particularly important as they directly reflect human physiological states, driving progress in areas such as disease diagnosis, patient monitoring, and treatment outcome prediction. Recent studies have explored various tasks, such as reconstruction Li et al. (2023a) and imputation Lim & Zohren (2021), to address issues of data incompleteness and noise. For example, ECG reconstruction aims to restore corrupted or missing signal segments caused by uncontrollable factors, ensuring the reliability and continuity of ECG data, which is essential for accurate clinical interpretation and decision-making.
Recently, score-based diffusion models have achieved state-of-the-art performance in time series reconstruction compared to traditional methods Wang et al. (2023); Alcaraz & Strodthoff (2022); Tashiro et al. (2021). The success in time series reconstruction can be attributed to their abilities to model complex data distributions, generating realistic outputs through iterative refinement, and automatic feature extraction in conjunction with inductive biases.
@@
Figure 1: The demonstration of the inconsistency in reconstruction quality among 12 different channels within a data sample and 4 different data samples. The data samples are the 12-channel ECG signals that are randomly selected from PTB-XL dataset and reconstructed by the SOTA method DeScoD-ECG Li et al. (2023a).
Hence, a single DNN model commonly struggles to achieve low reconstruction errors across all channels within the same sample and fails to maintain consistent performance across the same channels in different samples. Fig. 1 illustrates the limitations of the single model Li et al. (2023a) in handling the physiological time series dataset, as it struggles to maintain consistent performance across all channels within a single sample and fails to ensure stable performance across the same channel in different samples.
@@
In addition, the previous study Li et al. (2023a) has proven that performing multiple reconstructions of the physiological time series signals using the same model and averaging the results (K-shot Averaging) can effectively improve the accuracy of the reconstruction. However, this approach significantly increases the computational burden due to repeated inference, making it impractical for use in real-world medical applications. Based on this finding, and to avoid the significant computational burden, we innovatively leverage the nature of MoE to simultaneously generate multiple denoising factors for each channel at the output of the backbone, then utilize the fused denoising factor via router gates to reconstruct the signals in a single inference computation. Our observations show that our proposed fusion MoE shares a similar principle with the method that requires averaging multiple reconstructions, but our fusion MoE is much efficient. Furthermore, we theoretically prove that our fusion MoE outperforms averaging multiple reconstruction results in reducing the errors of reconstruction.
@@ 2.1 Diffusion Model
Diffusion models have recently gained significant attention in time series modeling, particularly for reconstruction, imputation, forecasting, and anomaly detection. Tashiro et al. (2023) introduced TimeDiff, which demonstrated superior performance in time series forecasting and classification. Kollovieh et al. (2023) proposed TSDiff, integrating self-guidance mechanisms to refine generative outputs, making it applicable to both unconditional and conditional time series synthesis. A broader survey by Lin et al. (2024) categorized diffusion-based approaches in time series and spatio-temporal data.
However, despite the rapid advancements in general time series modeling, diffusion-based approaches have not been extensively explored in the physiological time series domain. Some recent works have applied diffusion models for EHR data generation Li et al. (2023b) and physiological signal imputation Wang et al. (2022). Additionally, Wang et al. (2024) investigated their use in multi-modal medical applications, including physiological signals and medical imaging. Furthermore, diffusion-based models have been employed for time series anomaly detection Shen et al. (2023) and denoising in ICU patient monitoring Zhang et al. (2023). Despite these efforts, the development of diffusion models specifically tailored for physiological time series reconstruction remains limited.
@@ 2.2 Mixture of Experts
Despite these advancements, the application of MoE models in the medical time series domain remains underexplored. Notably, Lee & Hauskrecht (2022) proposed a residual MoE framework to adapt clinical sequences, highlighting the potential of MoE in this area. However, comprehensive studies leveraging MoE for medical time series reconstruction are scarce. Given the high heterogeneity across channels and significant intra-population variability in physiological time series data, in addition to the lack of dedicated MoE solutions, we propose a novel diffusion-based MoE framework specifically designed for reconstructing and imputing physiological time series signals. Our approach leverages diffusion-based generative modeling to enhance temporal consistency and adaptive expert selection to dynamically handle diverse signal characteristics, ensuring robust and accurate reconstructions across different patient populations. And the framework of the condition diffusion model used in this paper is shown in Fig 2, where $\bar{x}$ is the condition for each time step that guides the model to reconstruct signals.
@@ 3.1 RFAMoE and Bridge
The input to the RFAMoE block consists of NN feature maps, where $N=B\times C$. Here, $B$ represents the batch size, and $C$ denotes the number of channels in each sample. $T$ denotes the time length of the time series signals, and $L$ represents the number of channels in the feature maps. The RFAMoE block utilizes a combination of multi-scale convolution layers to extract features from channels with different receptive fields. The dimension NN encompasses all channels of all samples. When using the multi-scale convolutional expert group to process NN feature maps, each channel within the feature maps can independently select the most suitable kernel size, rather than enforcing a uniform kernel size with the same receptive field across channels at the same position within the NN feature maps Wang et al. (2025); Li et al. (2023a); Romero et al. (2021), which particularly disadvantageous for medical time series data that are highly sensitive to receptive fields. $k$ is the number of convolutional layers activated by each channel, referred to as top-$k$ policy. These extracted features are then weighted and aggregated based on weights assigned by the router, resulting in the final feature representation for each channel. In our method, we found that top-1 activation serves as the optimal candidate activation strategy for our datasets.
Following the module of multi-scale convolution expert layers, the instance normalization is applied to each channel, which facilitating convergence and maintain the statistical independence of features, ensuring they are not affected by inter-channel interference. The feature map after instance normalization is divided into two parts. One part undergoes a GELU activation, followed by element-wise multiplication with the other part Ma et al. (2024), explicitly increasing the network width. At the end of the RFAMoE block, the feature map $\mathbf{X}\in\mathbb{R}^{N\times T\times L}$ is reshaped to $\mathbf{X}\in\mathbb{R}^{B\times(C\times L)\times T}$, then the information interaction is performed across all channels along the batch dimension. The kernel size of the convolutional layer used for channel information fusion is set to 1, ensuring the most effective fusion of inter-channel information while preserving the long-sequence feature paradigm already established within each channel Donghao & Xue (2024). The final result of the information interaction is combined with the input through a residual connection to produce the output of the block.
@@ 3.2 Fusion MoE
Inspired by the K-shot averaging proposed by DeScoD-ECG Li et al. (2023a), and to avoid the substantial computational cost brought by K-shot averaging, we employ a fused MoE approach, enabling the network to generate a fused denoise factor in a single forward pass. Then reconstruct the signal using the fused denoise factor. In the Fusion MoE, the router computes gate weights for NN feature maps corresponding to $K$ expert convolution layers, each with a kernel size of 1, an input channel size of $L$, and an output channel size of 1. These gate weights are then used to fuse the weights and biases of the $K$ convolution layers into a single convolution layer. The fused convolution layer subsequently outputs the results for the corresponding feature map. The proposed framework of Fusion MoE is presented in Algorithm 1.
@@ 3.3 Network Architecture
We adopt the backbone structure from Li et al. (2023a), with the primary difference being the deployment of the Fusion MoE module at the end of the network and RFAMoE block.
@@ 4 Experiments
Experimental setup and baselines. We validate the superiority of our method by comparing it against three mainstream SOTA condition-based diffusion models on imputing medical time series data (SSSDS4 Alcaraz & Strodthoff (2022); Diffwave Kong et al. (2020), DeSco-ECG Li et al. (2023a)) for imputing ECG and PSG signals. In our experiments, both our method and DeSco-ECG Li et al. (2023a) were configured with a time step of 40 and a batch size of 6, while the number of training epochs was set to 120 for both our method and DeSco-ECG. We set SSSDS4 and Diffwave as their default settings, with a time step of 200, 100 training epochs, and a batch size of 128.
All reported times in this paper represent the inference time required to reconstruct a single complete sample on a single Nvidia A6000 GPU. In our method, the parameter $L$ is set to 160. In the Fusion MoE module, we note that increasing $K$ beyond 16 yields diminishing returns, as further discussed in the D.2 of Appendix D. The number of experts $K$ is set to 16, utilizing a top-1616 activation strategy, while RFAMoE employs a total of 15 convolution layer experts, utilizing the top-1 activation strategy. We compare our method with baseline methods using three similarity metrics (PRD, SSD, MAD) that quantify the quality of ECG and PSG signals reconstruction. Our accuracy results reported are the mean accuracy over 3 runs using different random seeds. In addition, we also statistic each model’s parameters, FLOPs, and reconstruction steps to provide a more comprehensive comparison.
@@ B.1 Comparison With MoE Methods
To complement the experiment results, we further provide additional comparisons with two open sourced MoE based methods (Time-MoE Shi et al. (2024) and Residual MoE Lee & Hauskrecht (2022)) on PTB imputation task. The results shown in Table  B.1 show that our method significantly exceeds the Time-MoE Shi et al. (2024) and Residual MoE Lee & Hauskrecht (2022) on PRD and SSD metrics. In addition, our method is also leading in the MAD metric, compared with Time-MoE and Residual MoE.

### v4

# Advancing time series completion via RFAMoE and MDFF

@@ 1 Introduction
Medical time series data have gained increasing attention with the advancement of deep neural networks. Among these, physiological time series signals, such as Electrocardiogram (ECG), Electroencephalogram (EEG), and Polysomnography (PSG), are particularly important as they directly reflect human physiological states, driving progress in areas such as disease diagnosis, patient monitoring, and treatment outcome prediction. Recent studies have explored various tasks, such as reconstruction Li et al. (2023) and imputation Lim and Zohren (2021), to address issues of data incompleteness and noise. For example, ECG reconstruction aims to restore corrupted or missing signal segments caused by uncontrollable factors, ensuring the reliability and continuity of ECG data, which is essential for accurate clinical interpretation and decision-making.
Recently, score-based diffusion models have achieved state-of-the-art performance in time series reconstruction compared to traditional methods Wang et al. (2023); Alcaraz and Strodthoff (2022); Tashiro et al. (2021). The success in time series reconstruction can be attributed to their abilities to model complex data distributions, generating realistic outputs through iterative refinement, and automatic feature extraction in conjunction with inductive biases.
@@
Figure 1: The demonstration of the inconsistency in reconstruction quality among 12 different channels within a data sample and 4 different data samples. The data samples are the 12-channel ECG signals that are randomly selected from PTB-XL dataset and reconstructed by the SOTA method DeScoD-ECG Li et al. (2023).
Hence, a single DNN model commonly struggles to achieve low reconstruction errors across all channels within the same sample and fails to maintain consistent performance across the same channels in different samples. Fig. 1 illustrates the limitations of the single model Li et al. (2023) in handling the physiological time series dataset, as it struggles to maintain consistent performance across all channels within a single sample and fails to ensure stable performance across the same channel in different samples.
@@
In addition, the previous study Li et al. (2023) has proven that performing multiple reconstructions of the physiological time series signals using the same model and averaging the results (K-shot Averaging) can effectively improve the accuracy of the reconstruction. However, this approach significantly increases the computational burden due to repeated inference, making it impractical for use in real-world medical applications. Based on this finding, and to avoid the significant computational burden, we innovatively leverage the nature of MoE to simultaneously generate multiple denoising factors for each channel at the output of the backbone, then utilize the fused denoising factor via router gates to reconstruct the signals in a single inference computation. Our observations show that our proposed fusion MoE shares a similar principle with the method that requires averaging multiple reconstructions, but our fusion MoE is much efficient. Furthermore, we theoretically prove that our fusion MoE outperforms averaging multiple reconstruction results in reducing the errors of reconstruction.
@@ 2.1 Diffusion Model
Diffusion models have recently gained significant attention in time series modeling, particularly for reconstruction, imputation, forecasting, and anomaly detection. Tashiro and others (2023) introduced TimeDiff, which demonstrated superior performance in time series forecasting and classification. Kollovieh et al. (2023) proposed TSDiff, integrating self-guidance mechanisms to refine generative outputs, making it applicable to both unconditional and conditional time series synthesis. A broader survey by Lin and others (2024) categorized diffusion-based approaches in time series and spatio-temporal data.
However, despite the rapid advancements in general time series modeling, diffusion-based approaches have not been extensively explored in the physiological time series domain. Some recent works have applied diffusion models for EHR data generation Li and others (2023) and physiological signal imputation Wang and others (2022). Additionally, Wang and others (2024) investigated their use in multi-modal medical applications, including physiological signals and medical imaging. Furthermore, diffusion-based models have been employed for time series anomaly detection Shen and others (2023) and denoising in ICU patient monitoring Zhang and others (2023). Despite these efforts, the development of diffusion models specifically tailored for physiological time series reconstruction remains limited.
@@ 2.2 Mixture of Experts
Despite these advancements, the application of MoE models in the medical time series domain remains underexplored. Notably, Lee and Hauskrecht (2022) proposed a residual MoE framework to adapt clinical sequences, highlighting the potential of MoE in this area. However, comprehensive studies leveraging MoE for medical time series reconstruction are scarce. Given the high heterogeneity across channels and significant intra-population variability in physiological time series data, in addition to the lack of dedicated MoE solutions, we propose a novel diffusion-based MoE framework specifically designed for reconstructing and imputing physiological time series signals. Our approach leverages diffusion-based generative modeling to enhance temporal consistency and adaptive expert selection to dynamically handle diverse signal characteristics, ensuring robust and accurate reconstructions across different patient populations. And the framework of the condition diffusion model used in this paper is shown in Fig 2, where $\bar{x}$ is the condition for each time step that guides the model to reconstruct signals.
@@ 3.1 RFAMoE and Bridge
The input to the RFAMoE block consists of $N$ feature maps, where $N=B\times C$. Here, $B$ represents the batch size, and $C$ denotes the number of channels in each sample. $T$ denotes the time length of the time series signals, and $L$ represents the number of channels in the feature maps. The RFAMoE block utilizes a combination of multi-scale convolution layers to extract features from channels with different receptive fields. The dimension $N$ encompasses all channels of all samples. When using the multi-scale convolutional expert group to process $N$ feature maps, each channel within the feature maps can independently select the most suitable kernel size, rather than enforcing a uniform kernel size with the same receptive field across channels at the same position within the $N$ feature maps Wang et al. (2025); Li et al. (2023); Romero et al. (2021), which particularly disadvantageous for medical time series data that are highly sensitive to receptive fields. $k$ is the number of convolutional layers activated by each channel, referred to as top-$k$ policy. These extracted features are then weighted and aggregated based on weights assigned by the router, resulting in the final feature representation for each channel. In our method, we found that top-1 activation serves as the optimal candidate activation strategy for our datasets.
Following the module of multi-scale convolution expert layers, the instance normalization is applied to each channel, which facilitating convergence and maintain the statistical independence of features, ensuring they are not affected by inter-channel interference. The feature map after instance normalization is divided into two parts. One part undergoes a GELU activation, followed by element-wise multiplication with the other part Ma et al. (2024), explicitly increasing the network width. At the end of the RFAMoE block, the feature map $\mathbf{X}\in\mathbb{R}^{N\times T\times L}$ is reshaped to $\mathbf{X}\in\mathbb{R}^{B\times(C\times L)\times T}$, then the information interaction is performed across all channels along the batch dimension. The kernel size of the convolutional layer used for channel information fusion is set to 1, ensuring the most effective fusion of inter-channel information while preserving the long-sequence feature paradigm already established within each channel Donghao and Xue (2024). The final result of the information interaction is combined with the input through a residual connection to produce the output of the block.
@@ 3.2 Fusion MoE
Inspired by the K-shot averaging proposed by DeScoD-ECG Li et al. (2023), and to avoid the substantial computational cost brought by K-shot averaging, we employ a fused MoE approach, enabling the network to generate a fused denoise factor in a single forward pass. Then reconstruct the signal using the fused denoise factor. In the Fusion MoE, the router computes gate weights for $N$ feature maps corresponding to $K$ expert convolution layers, each with a kernel size of 1, an input channel size of $L$, and an output channel size of 1. These gate weights are then used to fuse the weights and biases of the $K$ convolution layers into a single convolution layer. The fused convolution layer subsequently outputs the results for the corresponding feature map. The proposed framework of Fusion MoE is presented in Algorithm 1.
@@ 3.3 Network Architecture
We adopt the backbone structure from Li et al. (2023), with the primary difference being the deployment of the Fusion MoE module at the end of the network and RFAMoE block.
@@ 4 Experiments
Experimental setup and baselines. We validate the superiority of our method by comparing it against three mainstream SOTA condition-based diffusion models on imputing medical time series data (SSSDS4 Alcaraz and Strodthoff (2022); Diffwave Kong et al. (2020), DeSco-ECG Li et al. (2023)) for imputing ECG and PSG signals. In our experiments, both our method and DeSco-ECG Li et al. (2023) were configured with a time step of 40 and a batch size of 6, while the number of training epochs was set to 120 for both our method and DeSco-ECG. We set SSSDS4 and Diffwave as their default settings, with a time step of 200, 100 training epochs, and a batch size of 128.
All reported times in this paper represent the inference time required to reconstruct a single complete sample on a single Nvidia A6000 GPU. In our method, the parameter $L$ is set to 160. In the Fusion MoE module, we note that increasing $K$ beyond 16 yields diminishing returns, as further discussed in the D.2 of Appendix D. The number of experts $K$ is set to 16, utilizing a top-$16$ activation strategy, while RFAMoE employs a total of 15 convolution layer experts, utilizing the top-1 activation strategy. We compare our method with baseline methods using three similarity metrics (PRD, SSD, MAD) that quantify the quality of ECG and PSG signals reconstruction. Our accuracy results reported are the mean accuracy over 3 runs using different random seeds. In addition, we also statistic each model’s parameters, FLOPs, and reconstruction steps to provide a more comprehensive comparison.
@@ B.1 Comparison With MoE Methods
To complement the experiment results, we further provide additional comparisons with two open sourced MoE based methods (Time-MoE Shi et al. (2024) and Residual MoE Lee and Hauskrecht (2022)) on PTB imputation task. The results shown in Table  B.1 show that our method significantly exceeds the Time-MoE Shi et al. (2024) and Residual MoE Lee and Hauskrecht (2022) on PRD and SSD metrics. In addition, our method is also leading in the MAD metric, compared with Time-MoE and Residual MoE.